Grade 12th PassPhysical Chemistry

In the equation

N2 + H2 ---> NH3

i hav to find the limiting reagent by using EQUIVALENCE CONCEPT..! plz help (i m confused on how to find the gram equivalent of N2)

Nikhil Arora
15 Years agoGrade 12th Pass

1 Answer

RAJORSHI PAUL
15 Years ago

N2 + 3H2 ----> 2NH3

N2 + 6e -------> 2N (-3)

So each molecule of N2 is involved in 6e transfer. so eq wt of N2 is (mol wt.)/6 = 28/6=14/3=4.67